Ive got to solve these problems using Quadratic Equation, but obviously I cant. Help me out, thanks!
a) The product of the second and third consecutive integers is 7 more than the first. Find the numbers.
b) If the area of the triangle is 240 square meters and one leg is 8 more than the other leg, find the (sorry, n/a. please fill in)?

Answer by edjones(8007)

Let the numbers be x, x+1 and x+2.
(x+1)(x+2)=x+7
x^2+3x+2=x+7
x^2+2x-5=0
There is no answer because x is not an integer.
.
Ed